HERRERO’S APPROXIMATION PROBLEM FOR QUASIDIAGONAL OPERATORS

NATHANIAL P. BROWN

Abstract. Let T be a quasidiagonal operator on a separable Hilbert space. It is shown that there exists a sequence of operators {T_{n}} such that dim C^{∗}(T_{n}) < ∞ and ∥T − T_{n}∥ → 0 if and only if C^{∗}(T) is exact.